OpenStudy (anonymous):

The area of a regular octagon is 35 cm2. What is the area of a regular octagon with sides five times as long? A. 625 cm2 B. 875 cm2 C. 175 cm2 D. 245 cm2

OpenStudy (anonymous):

do i mulitply 35cm^2 * 5?

OpenStudy (anonymous):

Bc i got C last time and it was incorrect :/

OpenStudy (raden):

area of a octagon = 2*(1+sqrt(2))*S^2, with s=side so, we have 35 = 2*(1+sqrt(2))*S^2 or s^2 = 35/[(1+sqrt(2)] now, to find area of a octagon if the sides be 5 times before area = 2*(1+sqrt(2))*(5S)^2 = 2*(1+sqrt(2))*25S^2, subtitute s^2 to here = 2*(1+sqrt(2))*25*35/[(1+sqrt(2)] = 2*25*35 = 1750 lol, no answer in options

OpenStudy (anonymous):

oh well , i guess i'll just take my best guess :) lol thanks

OpenStudy (raden):

wait... looks i have mistaken above 35 = 2*(1+sqrt(2))*S^2 or s^2 = 35/(2*[(1+sqrt(2)]) the new area = 2*(1+sqrt(2))*(5S)^2 = 2*(1+sqrt(2))*25*S^2 = 2*(1+sqrt(2))*25* 35/(2*[(1+sqrt(2)]) = 25*35 = 875

OpenStudy (anonymous):

ohh okay understand ! unstead of multiplying 2*25*35 , you just multiply 25*35 , thanks so much!

OpenStudy (raden):

i was leave 2's the first, forget 2/2 = 1 sorry.. :)

OpenStudy (raden):

welcome